# Data Imputation by Pursuing Better Classification: A Supervised Kernel-Based Method

### v2

@@ 1 Introduction
The presence of missing data poses significant challenges in machine learning and data analysis [1, 2, 3, 4]. In many real-world application scenarios, ensuring that all entries in the data are complete is a difficult task, and this can be attributed to various reasons. Firstly, it may occur due to limitations in data collection, such as time constraints or limited resources, resulting in certain observations being omitted. Secondly, data can be intentionally designed to have missing values, for instance, in survey studies where participants may choose not to answer certain questions, thereby introducing missing values into the data set [5]. By completing missing data, researchers can uncover the underlying structure and relationships within the data set, optimizing the utilization of the complete data set to enhance the performance and efficiency of subsequent analyses. Therefore, the significance of filling in missing data becomes self-evident.
The existing data imputation methods primarily focus on the relationships among features. For example, mean imputation (MI, [6]) considers the mean value of the feature, while other methods consider the low-rank property of the imputed matrix [7, 8, 9, 10] or incorporate similarity information [11, 12, 13]. Some methods also utilize label information, with some performing different operations on data from different classes [14, 15], while others treat labels as additional features and then complete the matrix [16]. However, these utilizations of label information in imputation models lack flexibility and are insufficient to depict the complex relationship between data and labels. Considering a fundamental observation that data are helpful for distinguishing the labels, we can derive an imputation criterion:
@@
In a given sequence of triplets $\{(\boldsymbol{x}_{i},\boldsymbol{o}_{i},y_{i})\}_{i=1}^{N}$, where $\boldsymbol{x}_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\mathbb{R}$, we typically use the set of observable features $\boldsymbol{o}_{i}\subseteq 2^{\{1,\cdots,d\}}$ and missing features $\ast$ to represent the missing data $\boldsymbol{x}_{\boldsymbol{o}_{i}}^{i}\in(\mathbb{R}\cup\{\ast\})^{d}$. This sequence is then consolidated into an incomplete data set $\mathcal{D}=\{\boldsymbol{x}_{\boldsymbol{o}_{i}}^{i},y_{i}\}_{i=1}^{N}$. The task now becomes to find suitable missing values that aid in classification. For kernel-based methods widely applied in machine learning, such as support vector machine (SVM, [17, 18, 19, 20]), a significant characteristic is that classifiers can be explicitly written as functions of the relationship between data and labels. The relationship between data will be defined by a kernel function, and different kernel functions provide distinct measures of similarity. Specifically, a kernel matrix or Gram matrix, denoted as $\boldsymbol{K}\in\mathbb{R}^{N\times N}$, is constructed, where $K_{i,j}=k(\boldsymbol{x}_{i},\boldsymbol{x}_{j})$ represents the similarity between data $\boldsymbol{x}_{i}$ and $\boldsymbol{x}_{j}$ using a proper positive definite kernel function $k(\cdot,\cdot):\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$. A classifier is then trained based on this kernel matrix $\boldsymbol{K}$. When handling missing data, represented as $\boldsymbol{x}_{\boldsymbol{o}_{i}}^{i}$, the value of $K_{i,\cdot}\in\mathbb{R}^{N}$ is unknown. Thus, it is necessary to find $\tilde{K}_{i,j}$ for data with missing features in order to establish appropriate relationships with other data. In SVM, this process can be seamlessly integrated with supervised information, leading to enhanced prediction accuracy in subsequent classification tasks. By considering prior knowledge of missing values and imposing constraints on the adjustment of the kernel matrix, we can effectively incorporate this information into the modeling process. The next step is to recover the missing values from the imputed matrix $\tilde{\boldsymbol{K}}$. This step is relatively easy, particularly in the traditional scenario where $N\gg d$. Because for each data, there is $N-1$ available supervised information stored in the kernel matrix to guide the imputation of $d$ or fewer elements. In this paper, we propose a two-stage learning framework for data imputation: I) the kernel matrix is completed by pursuing high classification accuracy; II) the missing features are reconstructed to approximate the optimized kernel matrix. With the proposed method, we can find good estimates for missing values that lead to better classification performance in the subsequent task. Therefore, when faced with situations depicted in Figure 1, we are able to obtain the imputation results indicated by the red crosses and circles.
@@
In the first stage, specifically, we begin by computing the initial kernel matrix based on the observed features. This matrix is then further modified through a process referred to as kernel matrix imputation throughout this paper. Recent studies [21, 22] have demonstrated that integrating kernel learning and classifier optimization tasks, and alternating between them, can improve classification performance and naturally lead to a fine-tuned kernel matrix. The effectiveness of such joint learning ideas has recently been also confirmed in neural network-based methods as well [23]. Inspired by these ideas, we adopt a similar approach by combining the tasks of kernel matrix imputation and classifier training into a maximum-minimum optimization problem. However, for such joint optimization objectives, we need to ensure that the kernel matrix is optimized to achieve high classification accuracy while also making the classifier sufficiently robust. Otherwise, the flexibility of modifying the similarity relationships between data can make the classifier susceptible to overfitting. To prevent this situation, we introduce an additional perturbation matrix during the optimization of the kernel matrix. In the second stage, we utilize the block coordinate descent (BCD, [24]) method to solve a non-convex problem. Although finding the global optimal solution is challenging, our numerical experiments demonstrate that when the data size is much larger than the feature dimension, the abundant supervisory information from the kernel matrix effectively guides the imputation process. As a result, we can still achieve highly accurate solutions. The main contributions of this paper are summarized as follows:
@@
  • To fully leverage label information, we propose a novel two-stage data imputation framework. This framework optimizes the similarity relationships between data by utilizing supervision information and guides the data imputation process accordingly.
@@ 2 Related Work
Filling data before classification. A large portion of the research focuses on filling in the data before using it in the subsequent classification task. For instance, Ghahramani & Jordan. [25] introduced a framework based on maximum likelihood density estimation for coping with missing data, utilizing mixture models and the expectation-maximization principle. And Bhattacharyya et al. [26] modeled the missing values as Gaussian random variables and estimated their mean and covariance from the observed data and labels, which were then used in a modified SVM. Later, Williams et al. [27] used a Gaussian mixture model to estimate the conditional probability of missing values given the observed data and imputed them using expectations. Additionally, some work estimated complete data by computing the corresponding marginal distributions of the missing data [14] or assumed a low-rank subspace for the data and treated labels as an additional feature to handle missing values [16]. In recent years, Rabin. [28] proposed a kernel-based multi-scale data imputation method, and López. [29] utilized the known matrix structure in terms of subspaces to assist in matrix imputation. But these methods separate the imputation process from the classification task, which may not necessarily lead to improved performance with their imputation outcomes.
@@
Filling the kernel matrix before classification. Another research direction focuses on kernel-based machine learning algorithms. In this approach, classification or regression tasks only require the kernel matrix based on the training data. Therefore, methods have been explored to compute kernel function values between data with missing values instead of directly imputing the missing data. Hazan et al. [30] proposed a parameterized kernel function based on the low-rank assumption of data, which allows for input vectors with missing values. However, when computing similarities, only the dimensions present in both data are considered, and the rest are discarded. There is also work that modeled the squared distance between two missing data as a random variable following a gamma distribution and computed the expectation of the Gaussian kernel function under missing data [31]. Similarly, Śmieja et al. [13] modeled possible outcomes of missing values with the data distribution and computed the expectation of similarity between two missing data. Still, these methods have limitations in flexibly utilizing the supervision information from labels during the imputation process, which consequently restricts their performance in practical tasks.
@@
Direct classification of incomplete data. Researchers have also explored methods for directly classifying missing data. In the study by Pelckmans et al. [32], they defined a modified risk to address the uncertainty in prediction results caused by missing data. And in [33], they maximized the margin of the subspace correlated with the observed data. However, when computing kernel values, they encounter similar issues of incomplete information utilization as in [30]. In a different approach, Dekel et al. [34] treated missing values as a specific type of noise and developed a linear programming problem that resembled SVM formulation in order to learn a robust binary classifier. And some works optimize the nonlinear classifier while seeking the linear subspace where the data might lie [9, 10]. Due to the necessity of predefining the dimension of the subspace in these methods, they may fail to fully capture the intrinsic structure of the data, thus constraining its flexibility in practical applications. In addition to the aforementioned kernel-based methods, there are also numerous works in the field of neural networks that focus on classifying missing data. Recently, Śmieja et al. [35] replaced the responses of neurons in the first hidden layer with their expected values, thereby minimizing the extent of modifications when adapting to various networks. Later, a neural network architecture called NeuMiss was introduced [36]. It utilized a Neumann-series approximation of the optimal predictor and demonstrated robustness against various missing data mechanisms. However, as it essentially learns a linear model, its capability may be limited in complex scenarios. It is worth noting that neural network-based methods typically do not perform data imputation. On the one hand, training neural networks typically requires a large amount of data, making it difficult to impute missing values for small-scale data. And in fact, imputing small-scale data is even more critical and meaningful. This is because they inherently provide limited information, and imputation can be used to uncover the underlying structure. On the other hand, it may be challenging to control the outcome of data imputation due to the high degree of freedom during network optimization. Furthermore, apart from explorations in algorithmic applications, there are also dedicated efforts focused on theoretical derivations related to missing data. For example, Bullins et al. [37] provided the information-theoretic upper and lower bounds of precision limits for vanilla SVM in the context of learning with missing data. Josse et al. [38] proved that a predictor designed for complete observations can achieve optimal predictions on incomplete data by utilizing multiple imputation.
@@ 3.1 Preliminaries
[…]
We begin by introducing the vanilla SVM and then demonstrate how to fill in missing data to achieve improved classification performance. In the hard-margin SVM, the objective is to find a hyperplane $\boldsymbol{w}^{\top}\boldsymbol{x}+b=0$ that maximally separates the data $\boldsymbol{x}\in\mathbb{R}^{d}$ of different classes with zero training errors. When the data is not linearly separable, the soft-margin SVM is proposed, which allows for some training errors by introducing slack variables denoted as $\xi_{i}\geq 0$ for each training sample. Meanwhile, in order to find a more flexible decision boundary, a nonlinear mapping $\phi(\cdot)$ from the $\mathbb{R}^{d}$ to a reproducing kernel Hilbert space $\mathcal{H}_{k}$ is introduced, yielding the decision function $f(\boldsymbol{x})=\mathrm{sign}(\boldsymbol{w}^{\top}\phi(\boldsymbol{x})+b)$. With the completed data, the optimization problem can be formulated as
@@
$\mathrm{s.t.}$ $y_{i}(\boldsymbol{w}^{\top}\phi(\boldsymbol{x}_{i})+b)\geq 1-\xi_{i},$
$\xi_{i}\geq 0,\ \forall i\in[N],$
@@
[…]
where $\boldsymbol{y}=[y_{1};y_{2}\ ;\ldots;y_{N}]\in\mathbb{R}^{N}$ and each entry $K_{i,j}=\langle\phi(\boldsymbol{x}_{i}),\phi(\boldsymbol{x}_{j})\rangle_{\mathcal{H}_{k}}$ of the kernel matrix $\boldsymbol{K}\in\mathbb{R}^{N\times N}$ represents the similarity between two data. Through a technique known as the kernel trick, this similarity can be computed using a predefined positive definite kernel function $k(\boldsymbol{x}_{i},\boldsymbol{x}_{j}):\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$, allowing us to calculate it without knowing the explicit expressions of $\phi$. And the resulting kernel matrix $\boldsymbol{K}$ is guaranteed to be PSD.
@@ 3.2 Stage I: Kernel Matrix Imputation with SVM
In this stage, our goal is to find a new kernel matrix $\tilde{\boldsymbol{K}}$ that further reduces the objective function value in (2), aiming to improve the classification accuracy of the trained classifier. We propose a novel algorithm that integrates kernel matrix imputation with the classification task. Initially, we compute the observed kernel matrix $\boldsymbol{K}_{o}$, and then optimize an adjustment matrix $\boldsymbol{K}_{\Delta}$ with the same dimension to derive $\tilde{\boldsymbol{K}}=\boldsymbol{K}_{o}\odot\boldsymbol{K}_{\Delta}$ as the optimized kernel matrix. This approach distinguishes itself from previous algorithms by effectively utilizing each observed feature value. Considering the Gaussian kernel function $k_{\gamma}(\boldsymbol{x},\boldsymbol{y})=\exp(-\gamma\|\boldsymbol{x}-\boldsymbol{y}\|_{2}^{2})$ as an example and recalling the definition of the incomplete data set $\mathcal{D}=\{\boldsymbol{x}_{\boldsymbol{o}_{i}}^{i},y_{i}\}_{i=1}^{N}$, we can decompose $K_{i,j}$ into an observed part and unknown parts:
@@
Recall that $\ast$ represents an unknown real number. Methods in [33, 30] only consider the similarity between the intersecting observed features of two data, i.e., the part related to $p\in\boldsymbol{o}_{i}\cap\boldsymbol{o}_{j}$. In doing so, they disregard the features that are observed in one data but missing in the other, namely $x_{p}^{i}$ $(p\in\boldsymbol{o}_{i}\backslash\boldsymbol{o}_{j})$ and $x_{p}^{j}$ $(p\in\boldsymbol{o}_{j}\backslash\boldsymbol{o}_{i})$. In contrast, our algorithm calculates $\boldsymbol{K}_{o}$ using the first term in the last line of (3) and implicitly incorporates the remaining terms by constraining the potential ranges for $\boldsymbol{K}_{\Delta}$. Assuming $\boldsymbol{x}$ is min-max normalized within the range of $[0,1]$, the ranges of the second, third, and fourth terms in (3) are denoted as $[0,\sum_{p\in\boldsymbol{o}_{i}\backslash\boldsymbol{o}_{j}}\max\{(x_{p}^{i})^{2},(1-x_{p}^{i})^{2}\}]$, $[0,\sum_{p\in\boldsymbol{o}_{j}\backslash\boldsymbol{o}_{i}}\max\{(x_{p}^{j})^{2},(1-x_{p}^{j})^{2}\}]$, and $[0,d-|\boldsymbol{o}_{i}\cup\boldsymbol{o}_{j}|]$, respectively. By computing the range of each entry $K_{i,j}=(K_{o})_{i,j}\cdot(K_{\Delta})_{i,j}$, we can determine the feasible domain of the optimization variable $\boldsymbol{K}_{\Delta}$, denoted as $\boldsymbol{B}_{l}\preceq\boldsymbol{K}_{\Delta}\preceq\boldsymbol{B}_{u}$. In Stage II and the experimental sections, we will continue to use the Gaussian kernel as an example. However, it is important to emphasize that our algorithm can be applied to various kernel functions, as long as their computations can be decoupled for each feature. This decoupling can be expressed as
@@
where the range of the binary function $g(\cdot,\cdot)$ needs to be bounded when both inputs in the range $[0,1]$, and $f(\cdot)$ needs to be a monotonic function on the range of $\sum_{p=1}^{d}g\left(x_{i},y_{i}\right)$. Many commonly used kernels exhibit this property, including the polynomial kernel, Laplacian kernel, sigmoid kernel, $\chi^{2}$ kernel, and more. For a detailed derivation of the decoupling for these kernels, please refer to the Appendix A.
By leveraging the supervision information of the labels and performing alternating optimization between the kernel matrix and the classifier, we will obtain a kernel matrix that yields better classification performance. Meanwhile, due to the flexibility of this approach, we aim to ensure the robustness of the classifier concurrently, thus mitigating the risk of overfitting. Therefore, based on the concept of robust optimization [39, 40], we introduce a perturbation variable denoted as $\mathcal{E}$ in the space of $\mathbb{R}^{N\times N}$ to guarantee that the classifier performs well not only on $\tilde{\boldsymbol{K}}$ but also on all possible outcomes within a norm sphere surrounding $\tilde{\boldsymbol{K}}$. Finally, we formulate the following optimization problem for Stage I:
$\min_{\boldsymbol{K}_{\Delta}}\max_{\boldsymbol{\alpha},\mathcal{E}}$ $\boldsymbol{1}^{\top}\boldsymbol{\alpha}-\frac{1}{2}\boldsymbol{\alpha}^{\top}\mathrm{diag}(\boldsymbol{y})\left(\boldsymbol{K}_{o}\odot\boldsymbol{K}_{\Delta}\odot\mathcal{E}\right)\mathrm{diag}(\boldsymbol{y})\boldsymbol{\alpha}+\eta\|\boldsymbol{K}_{\Delta}-\boldsymbol{1}\boldsymbol{1}^{\top}\|_{\mathrm{F}}^{2}$ (5)
@@
[…]

### v3

@@ Abstract
journal: Pattern Recognition
[…]
[sjtu]organization=Institute of Image Processing and Pattern Recognition, Department of Automation, Shanghai Jiao Tong University, city=Shanghai, postcode=200240, country=China
@@ 1 Introduction
The presence of missing data poses significant challenges in machine learning and data analysis [1, 2, 3]. In many real-world application scenarios, ensuring that all entries in the data are complete is a difficult task, and this can be attributed to various reasons. Firstly, it may occur due to limitations in data collection, such as time constraints or limited resources, resulting in certain observations being omitted. Secondly, data can be intentionally designed to have missing values, for instance, in survey studies where participants may choose not to answer certain questions, thereby introducing missing values into the dataset. By completing missing data, researchers can uncover the underlying structure and relationships within the dataset, optimizing the utilization of the complete dataset to enhance the performance and efficiency of subsequent analyses. Therefore, the significance of filling in missing data becomes self-evident.
The existing data imputation methods primarily focus on the relationships among features. For example, mean imputation (MI, [4]) considers the mean value of the feature, while other methods consider the low-rank property of the imputed matrix [5, 6, 7] or incorporate similarity information [8, 9]. Some methods also utilize label information, with some performing different operations on data from different classes [10, 11], while others treat labels as additional features and then complete the matrix [12]. However, these utilizations of label information in imputation models lack flexibility and are insufficient to depict the complex relationship between data and labels. Considering a fundamental observation that data are helpful for distinguishing the labels, we can derive an imputation criterion:
@@
In a given sequence of triplets $\{(\boldsymbol{x}_{i},\boldsymbol{o}_{i},y_{i})\}_{i=1}^{N}$, where $\boldsymbol{x}_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\mathbb{R}$, we typically use the set of observable features $\boldsymbol{o}_{i}\subseteq 2^{\{1,\cdots,d\}}$ and missing features $\ast$ to represent the missing data $\boldsymbol{x}_{\boldsymbol{o}_{i}}^{i}\in(\mathbb{R}\cup\{\ast\})^{d}$. This sequence is then consolidated into an incomplete dataset $\mathcal{D}=\{\boldsymbol{x}_{\boldsymbol{o}_{i}}^{i},y_{i}\}_{i=1}^{N}$. The task now becomes to find suitable missing values that aid in classification. For kernel-based methods widely applied in machine learning, such as support vector machine (SVM, [13, 14, 15]), a significant characteristic is that classifiers can be explicitly written as functions of the relationship between data and labels. The relationship between data will be defined by a kernel function, and different kernel functions provide distinct measures of similarity. Specifically, a kernel matrix or Gram matrix, denoted as $\boldsymbol{K}\in\mathbb{R}^{N\times N}$, is constructed, where $K_{i,j}=k(\boldsymbol{x}_{i},\boldsymbol{x}_{j})$ represents the similarity between data $\boldsymbol{x}_{i}$ and $\boldsymbol{x}_{j}$ using a proper positive definite kernel function $k(\cdot,\cdot):\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$. A classifier is then trained based on this kernel matrix $\boldsymbol{K}$. When handling missing data, represented as $\boldsymbol{x}_{\boldsymbol{o}_{i}}^{i}$, the value of $K_{i,\cdot}\in\mathbb{R}^{N}$ is unknown. Thus, it is necessary to find $\tilde{K}_{i,j}$ for data with missing features in order to establish appropriate relationships with other data. In SVM, this process can be seamlessly integrated with supervised information, leading to enhanced prediction accuracy in subsequent classification tasks. By considering prior knowledge of missing values and imposing constraints on the adjustment of the kernel matrix, we can effectively incorporate this information into the modeling process. The next step is to recover the missing values from the imputed matrix $\tilde{\boldsymbol{K}}$. This step is relatively easy, particularly in the traditional scenario where $N\gg d$. Because for each data, there is N1N-1 available supervised information stored in the kernel matrix to guide the imputation of $d$ or fewer elements. In this paper, we propose a two-stage framework for data imputation: I) the kernel matrix is completed by pursuing high classification accuracy; II) the missing features are reconstructed to approximate the optimized kernel matrix. With the proposed method, we can find good estimates for missing values that lead to better classification performance in the subsequent task. Therefore, when faced with situations depicted in Figure 1, we are able to obtain the imputation results indicated by the red crosses and circles.
@@
In the first stage, we begin by computing the initial kernel matrix based on the observed features. This matrix is then further modified through a process referred to as kernel matrix imputation throughout this paper. Recent studies [16, 17] have demonstrated that integrating kernel learning and classifier optimization tasks, and alternating between them, can improve classification performance and naturally lead to a fine-tuned kernel matrix. The effectiveness of such joint learning ideas has recently been also confirmed in neural network-based methods as well [18]. Inspired by these ideas, we adopt a similar approach by combining the tasks of kernel matrix imputation and classifier training into a maximum-minimum optimization problem. However, for such joint optimization objectives, we need to ensure that the kernel matrix is optimized to achieve high classification accuracy while also making the classifier sufficiently robust. Otherwise, the flexibility of modifying the similarity relationships between data can make the classifier susceptible to overfitting. To prevent this situation, we introduce an additional perturbation matrix during the optimization of the kernel matrix. In the second stage, we utilize the block coordinate descent (BCD, [19]) method to solve a non-convex problem. Although finding the global optimal solution is challenging, our numerical experiments demonstrate that when the data size is much larger than the feature dimension, the abundant supervisory information from the kernel matrix effectively guides the imputation process. As a result, we can still achieve highly accurate solutions. The main contributions of this paper are summarized as follows:
  • 1.
@@
    To fully leverage label information, we propose a novel two-stage imputation framework. This framework optimizes the similarity relationships between data by utilizing supervision information and guides the imputation accordingly.
  • 2.
@@ 2 Related Work
Filling data before classification. A large portion of the research focuses on filling in the data before using it in the subsequent classification task. For instance, a framework based on maximum likelihood density estimation for coping with missing data was introduced, utilizing mixture models and the expectation-maximization principle [20]. A method that modeled the missing values as Gaussian random variables and estimated their mean and covariance from the observed data and labels, which were then used in a modified SVM, was proposed [21]. Later, a Gaussian mixture model was employed to estimate the conditional probability of missing values given the observed data and impute them using expectations [22]. Additionally, some work estimated complete data by computing the corresponding marginal distributions of the missing data [10] or assumed a low-rank subspace for the data and treated labels as an additional feature to handle missing values [12]. In recent years, a kernel-based multi-scale data imputation method was proposed [23], and the known matrix structure in terms of subspaces was utilized to assist in matrix imputation [24]. In addition, there are currently many deep learning-based imputation methods, such as those based on autoencoders [25], generative adversarial networks [26, 27], attention-based approaches [28], and methods leveraging diffusion models [29, 30]. But these methods separate the imputation process from the classification task, which may not necessarily lead to improved performance with their imputation outcomes.
@@
Filling the kernel matrix before classification. Another research direction focuses on kernel-based machine learning algorithms. In this approach, classification or regression tasks only require the kernel matrix based on the training data. Therefore, methods have been explored to compute kernel function values between data with missing values instead of directly imputing the missing data. A parameterized kernel function based on the low-rank assumption of data was proposed, which allows for input vectors with missing values [31]. However, when computing similarities, only the dimensions present in both data are considered, and the rest are discarded. There is also work that modeled the squared distance between two missing data as a random variable following a gamma distribution and computed the expectation of the Gaussian kernel function under missing data [32]. Similarly, possible outcomes of missing values were modeled with the data distribution, and the expectation of similarity between two missing data was computed [9]. Still, these methods have limitations in flexibly utilizing the supervision information from labels during the imputation process, which consequently restricts their performance in practical tasks.
@@
Direct classification of incomplete data. Researchers have also explored methods for directly classifying missing data. In one study, a modified risk was defined to address the uncertainty in prediction results caused by missing data [33]. Another approach maximized the margin of the subspace correlated with the observed data [34]. However, when computing kernel values, they encounter similar issues of incomplete information utilization as in [31]. In a different approach, missing values were treated as a specific type of noise, and a linear programming problem resembling SVM formulation was developed to learn a robust binary classifier [35]. And some works optimize the nonlinear classifier while seeking the linear subspace where the data might lie [36, 7]. Due to the necessity of predefining the dimension of the subspace in these methods, they may fail to fully capture the intrinsic structure of the data, thus constraining its flexibility in practical applications. In addition to the aforementioned kernel-based methods, numerous works in the field of neural networks focus on classifying missing data. For instance, the responses of neurons in the first hidden layer were replaced with their expected values to minimize the extent of modifications when adapting to various networks [37]. Later, a neural network architecture called NeuMiss was introduced [38]. It utilized a Neumann-series approximation of the optimal predictor and demonstrated robustness against various missing data mechanisms. However, as it essentially learns a linear model, its capability may be limited in complex scenarios. It is worth noting that training neural networks typically requires a large amount of data, making it difficult to impute missing values for small-scale data. And in fact, imputing small-scale data is even more critical and meaningful. This is because they inherently provide limited information, and imputation can be used to uncover the underlying structure. On the other hand, it may be challenging to control the outcome of data imputation due to the high degree of freedom during network optimization. Furthermore, apart from explorations in algorithmic applications, there are also dedicated efforts focused on theoretical derivations related to missing data. For example, the information-theoretic upper and lower bounds of precision limits for vanilla SVM in the context of learning with missing data were provided [39]. It was also proven that a predictor designed for complete observations can achieve optimal predictions on incomplete data by utilizing multiple imputation [40].
@@ 3.1 Preliminaries
Notations. The set of real numbers is written as $\mathbb{R}$. The set of integers from 1 to $N$ is written as $[N]$. The cardinality of the set $\mathcal{A}$ is written as $|\mathcal{A}|$. We take $a$, $\boldsymbol{a}$, and $\boldsymbol{A}$ to be a scalar, a vector, and a matrix, respectively. Let $\boldsymbol{0}$ and $\boldsymbol{1}$ denote vectors consisting of all zeros and all ones with the appropriate size. The inner product of two vectors in the given space is written as $\langle\cdot,\cdot\rangle$. We take $\mathrm{diag}(\boldsymbol{a})$ to be an operator that extends a vector to a diagonal matrix and $\mathrm{vec}(\boldsymbol{A})$ to be an operator that converts a matrix into a vector by stacking the columns of $\boldsymbol{A}$ on top of one another. The Frobenius norm is written as $\|\cdot\|_{\mathrm{F}}$. The set of symmetric matrices is denoted as $\mathcal{S}$, while the set of positive semi-definite (PSD) matrices is denoted as $\mathcal{S}_{+}$. The Hadamard product is written as $\odot$.
We introduce the vanilla SVM and then demonstrate how to fill in missing data to achieve improved classification performance. In the hard-margin SVM, the objective is to find a hyperplane $\boldsymbol{w}^{\top}\boldsymbol{x}+b=0$ that maximally separates the data $\boldsymbol{x}\in\mathbb{R}^{d}$ of different classes with zero training errors. When the data is not linearly separable, the soft-margin SVM is proposed, which allows for some training errors by introducing slack variables denoted as $\xi_{i}\geq 0$ for each training sample. Meanwhile, to find a more flexible decision boundary, a nonlinear mapping $\phi(\cdot)$ from the $\mathbb{R}^{d}$ to a reproducing kernel Hilbert space $\mathcal{H}_{k}$ is introduced, yielding the decision function $f(\boldsymbol{x})=\mathrm{sign}(\boldsymbol{w}^{\top}\phi(\boldsymbol{x})+b)$. With the completed data, the optimization problem can be formulated as
@@
$\mathrm{s.t.}$ $y_{i}(\boldsymbol{w}^{\top}\phi(\boldsymbol{x}_{i})+b)\geq 1-\xi_{i},\ \xi_{i}\geq 0,\ \forall i\in[N],$
@@
$\mathrm{s.t.}$ $\boldsymbol{y}^{\top}\boldsymbol{\alpha}=0,\boldsymbol{0}\leq\boldsymbol{\alpha}\leq C\boldsymbol{1},$
where $\boldsymbol{y}=[y_{1};y_{2}\ ;\ldots;y_{N}]\in\mathbb{R}^{N}$ and each entry $K_{i,j}=\langle\phi(\boldsymbol{x}_{i}),\phi(\boldsymbol{x}_{j})\rangle_{\mathcal{H}_{k}}$ of the kernel matrix $\boldsymbol{K}\in\mathbb{R}^{N\times N}$ represents the similarity between two data. Through a technique known as the kernel trick, this similarity can be computed using a predefined positive definite kernel function $k(\boldsymbol{x}_{i},\boldsymbol{x}_{j}):\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$, allowing us to calculate it without knowing the explicit expressions of $\phi$. And the resulting kernel matrix $\boldsymbol{K}$ is guaranteed to be PSD. Next, we will introduce our proposed two-stage data imputation method stage by stage. A schematic diagram of the method is shown in Figure 2.
Figure 2: Schematic diagram of the two-stage data imputation framework.
@@ 3.2 Stage I: Kernel Matrix Imputation with SVM
In this stage, our goal is to find a new kernel matrix $\tilde{\boldsymbol{K}}$ that further reduces the objective function value in problem (2), aiming to improve the classification accuracy of the trained classifier. We propose a novel algorithm that integrates kernel matrix imputation with the classification task. Initially, we compute the observed kernel matrix $\boldsymbol{K}_{o}$, and then optimize an adjustment matrix $\boldsymbol{K}_{\Delta}$ with the same dimension to derive $\tilde{\boldsymbol{K}}=\boldsymbol{K}_{o}\odot\boldsymbol{K}_{\Delta}$ as the optimized kernel matrix. This approach distinguishes itself from previous algorithms by effectively utilizing each observed feature value. Considering the Gaussian kernel function $k_{\gamma}(\boldsymbol{x},\boldsymbol{y})=\exp(-\gamma\|\boldsymbol{x}-\boldsymbol{y}\|_{2}^{2})$ as an example and recalling the definition of the incomplete dataset $\mathcal{D}=\{\boldsymbol{x}_{\boldsymbol{o}_{i}}^{i},y_{i}\}_{i=1}^{N}$, we can decompose $K_{i,j}$ into an observed part and unknown parts:
@@
Recall that $\ast$ represents an unknown real number. Methods in [34, 31] only consider the similarity between the intersecting observed features of two data, i.e., the part related to $p\in\boldsymbol{o}_{i}\cap\boldsymbol{o}_{j}$. In doing so, they disregard the features that are observed in one data but missing in the other, namely $x_{p}^{i}$ $(p\in\boldsymbol{o}_{i}\backslash\boldsymbol{o}_{j})$ and $x_{p}^{j}$ $(p\in\boldsymbol{o}_{j}\backslash\boldsymbol{o}_{i})$. In contrast, our algorithm calculates $\boldsymbol{K}_{o}$ using the first term in the last line of equation (3) and implicitly incorporates the remaining terms by constraining the potential ranges for $\boldsymbol{K}_{\Delta}$. Assuming $\boldsymbol{x}$ is min-max normalized within the range of $[0,1]$, the ranges of the second, third, and fourth terms in equation (3) are denoted as $0,\sum_{p\in\boldsymbol{o}_{i}\backslash\boldsymbol{o}_{j}}\max\{(x_{p}^{i})^{2},(1-x_{p}^{i})^{2}\}]$, $0,\sum_{p\in\boldsymbol{o}_{j}\backslash\boldsymbol{o}_{i}}\max\{(x_{p}^{j})^{2},(1-x_{p}^{j})^{2}\}]$, and $0,d-|\boldsymbol{o}_{i}\cup\boldsymbol{o}_{j}|]$, respectively. By computing the range of each entry $K_{i,j}=(K_{o})_{i,j}\cdot(K_{\Delta})_{i,j}$, we can determine the feasible domain of the optimization variable $\boldsymbol{K}_{\Delta}$, denoted as $\boldsymbol{B}_{l}\preceq\boldsymbol{K}_{\Delta}\preceq\boldsymbol{B}_{u}$. In Stage II and the experimental sections, we will continue to use the Gaussian kernel as an example. However, it is noted that our algorithm can be applied to various kernel functions, as long as their computations can be decoupled for each feature. This decoupling can be expressed as
@@
where the range of the binary function $g(\cdot,\cdot)$ needs to be bounded when both inputs in the range $[0,1]$, and $f(\cdot)$ needs to be a monotonic function on the range of $\sum_{p=1}^{d}g\left(x_{i},y_{i}\right)$. Many commonly used kernels exhibit this property, including the polynomial kernel, sigmoid kernel, $\chi^{2}$ kernel, and more. For a detailed derivation of the decoupling for these kernels, please refer to the A.
By leveraging the supervision information and performing alternating optimization between the kernel matrix and the classifier, we will obtain a kernel matrix that yields better classification performance. Meanwhile, due to the flexibility of this approach, we aim to ensure the robustness of the classifier concurrently, thus mitigating the risk of overfitting. Therefore, based on the concept of robust optimization [41, 42], we introduce a perturbation variable denoted as $\mathcal{E}\in\mathbb{R}^{N\times N}$ to guarantee that the classifier performs well not only on $\tilde{\boldsymbol{K}}$ but also on all possible outcomes within a norm sphere surrounding $\tilde{\boldsymbol{K}}$. Finally, we formulate the following optimization problem for Stage I:
$\min_{\boldsymbol{K}_{\Delta}}\max_{\boldsymbol{\alpha},\mathcal{E}}$ $\mathcal{L}=\boldsymbol{1}^{\top}\boldsymbol{\alpha}-\frac{1}{2}\boldsymbol{\alpha}^{\top}\mathrm{diag}(\boldsymbol{y})\left(\boldsymbol{K}_{o}\odot\boldsymbol{K}_{\Delta}\odot\mathcal{E}\right)\mathrm{diag}(\boldsymbol{y})\boldsymbol{\alpha}+\eta\|\boldsymbol{K}_{\Delta}-\boldsymbol{1}\boldsymbol{1}^{\top}\|_{\mathrm{F}}^{2}$ (5)
@@
$\boldsymbol{y}^{\top}\boldsymbol{\alpha}=0,\boldsymbol{0}\leq\boldsymbol{\alpha}\leq C\boldsymbol{1},$